OpenStudy (anonymous):

Marna wants to store 150 cubic inches of homemade pasta sauce in a food storage container. Should she choose the cylinder with a radius of 3 inches and a height of 5 inches, or the cube with a side length of 6 inches? Explain.

OpenStudy (mathstudent55):

You need to find the volumes of the two containers. \(\large {V_{cylinder} = \pi r^2 h} \) \(\large{V_{cube} = s^3 }\)